# S-IQA: IMAGE QUALITY ASSESSMENT WITH COMPRESSIVE SAMPLING

### v1

@@ Abstract
No-Reference Image Quality Assessment (IQA) aims at estimating image quality in accordance with subjective human perception. However, most existing NR-IQA methods focus on exploring increasingly complex networks or components to improve the final performance. Such practice imposes great limitations and complexity on IQA methods, especially when they are applied to high-resolution (HR) images in the real world. Actually, most images own high spatial redundancy, especially for those HR data. To further exploit the characteristic and alleviate the issue above, we propose a new framework for Image Quality Assessment with compressive Sampling (dubbed S-IQA), which consists of three components: (1) The Flexible Sampling Module (FSM) samples the image to obtain measurements at an arbitrary ratio. (2) Vision Transformer with the Adaptive Embedding Module (AEM) makes measurements of uniform size and extracts deep features (3) Dual Branch (DB) allocates weight for every patch and predicts the final quality score. Experiments show that our proposed S-IQA achieves state-of-the-art result on various datasets with less data usage.
Index Terms—  Image Quality Assessment, Compressive Sensing, Vision Transformer
1 Introduction
When we access an image from the Internet, it may have passed the following processing stages: captured from a camera, compressed when shared online, spread from people to people, and finally received. Any distortion can be introduced during these stages, which makes our final received image badly degraded from the original one and leaves us annoyed. Low-quality images can also cause deadly problems, especially in autonomous driving. Given these reasons, assessing the quality of images becomes particularly important in our daily lives.
Objective Image Quality Assessment (IQA) employs computational models to provide a quality score of an image from a human’s perspective. Based on the availability of the lossless reference image, objective quality metrics can be divided into full-reference (FR-IQA) and no-reference (NR-IQA) methods. NR-IQA methods are divided mainly into two categories: distortion-based and general-proposed methods. The former aims at predicting quality score towards a certain distortion type, leading to excellent performance but weak generalization. General-proposed methods deal with any kind of distortion but strongly depend on suitable feature extraction. However, as images of high resolution emerge, both methods meet severe computation burden. Actually, instead of the whole image, we might only need a small number of measurements of an image (perhaps 20% measurements) to achieve highly reliable score prediction. In implementation, we employ compressive sensing for sampling.
Compressive sensing (CS) is a novel method that breaks through the limitations of the Nyquist theorem on signal sampling. It demonstrates that a signal can be reconstructed from rather few measurements from the original signal. However, in many applications, such as IQA, the precise reconstructed image isn’t a primary concern. Still, we emphasize more on the final score, which means we can directly perform high-level tasks on the compressed domain, i.e., compressive learning.
The concept of compressive learning (CL) was first proposed by Calderbank et al.[1] and Davenport et al.[2] when they built the inference system directly using measurements without reconstruction. In fact, there are a lot of similarities between measurements and the original image. [3] makes some theoretical works and gradually finds some multidimensional properties of CL. A recent work [4] by Chong Mou et al. employs transformer as the backbone to cope with information loss through element-wise correlations and achieves amazing performance.
Fig. 1: The architecture of our proposed S-IQA. A distorted image is cropped into non-overlapping 16×16 sized patches. Then the patches are sampled by the FSM to get the compressed measurements. To further extract high-level features, we employ ViT. Moreover, to meet the input shape of transformer, we utilize an Adaptive Embedding Module (AEM) to reshape the measurements. Finally, a dual-branch structure is employed to get the score and weight of each patch and predict the quality score.
Inspired by this work, to make NR-IQA more data-efficient, we propose S-IQA. As shown in Fig.1, S-IQA consists of three parts: A Flexible Sampling Module (FSM), a Vision Transformer with the Adaptive Embedding Module (AEM), and a Dual Branch (DB) for final scoring. The original image is sampled in the flexible sampling module based on the input sampling ratio. Subsequently, the measurements are embedded by an adaptive embedding module and fed into the vision transformer to extract high-level features. Finally, a dual-branch structure is employed for the score prediction.
Our main contributions are as follows: (1) We apply compressive sensing to image quality assessment, enable sampling of images at arbitrary ratios, and achieve amazing performance with a small amount of data. (2) We employ an adaptive embedding module to handle the irregular shape caused by the aforementioned arbitrary sampling ratio. In this way, we can deal with kinds of ratios after training only once. (3) A dual-branch structure is proposed for quality score. In DB, we design scoring and weighting branches for the quality and weight of each patch. We assume that the scores of different parts of the image matter differently, and thus each patch should have its own score and weight.
Fig. 2: Random-crop removes pixels from every block despite the importance, leading to severe information loss.
2 RELATED WORK
2.1 Compressive Sensing
Existing CS methods can be classified into two classes: optimization-based and deep-learning-based methods. Traditional optimization-based methods mainly rely on sparsity priors to recover signals from measurements. Some methods attempt to solve it through L1 minimization, such as ISTA, AMP, and BP. There are also data-driven methods that are considered to be more effective than traditional ones, such as [5], [6]. However, the major issue with traditional methods is their extremely high computational complexity due to the iterative calculations. With the development of deep learning, Shi et al.[7] propose a network model called CSNet and show its promising performance. Moreover, some models that mimic traditional methods begin to emerge, such as ISTA-Net and OPINE-Net. Bin Chen et al.[8] use content-aware technique to guide the allocation of sampling ratio and subsequent reconstruction.
2.2 Image Quality Assessment
@@
Unlike FR-IQA, NR-IQA does not have access to reference images. Before the rise of deep learning, traditional methods were mainly divided into natural scene statistics (NSS) metrics and learning-based metrics. The assumption behind this is that the natural scene statistics (NSS) extracted from natural images tend to be highly regular, so distortions to an image would break these statistical regularities and make them detectable. With the success of deep learning and convolution neural networks (CNNs) in computer vision tasks, CNN-based methods have significantly outperformed traditional approaches in handling real-world distortions. Early deep learning NR-IQA methods employed stacked CNNs for feature extraction. [9] proposes a multi-task network where two sub-networks are trained for distortion identification and quality assessment. Hyper-IQA[10] leverages both low-level and high-level features and makes the latter redirect the former. Furthermore, [11] proposes a model that employs meta-learning to capture shared prior knowledge among different distortions.
Fig. 3: Process of Flexible Sampling Module (FSM). The input image is split into non-overlapping blocks and then sampled by the same matrix.
3 PROPOSED METHOD
@@
In this section, we elaborate on the details of our framework and integrated model. The whole architecture of our S-IQA is presented in Fig.1, which is mainly composed of a Flexible Sampling Module (FSM), a Vision Transformer with Adaptive Embedding Module (AEM) and a Dual Branch (DB).
Fig. 4: Illustration of Vision Transformer with Adaptive Embedding Module (AEM). The random-shape measurements are embedded into a unified dimension by AEM before fed into ViT for feature extraction.
3.1 Flexible Sampling Module
Among the recent works, there already exist some data-efficient methods for quality assessment, such as FAST-VQA[12] with random-crop. But cropping cannot guarantee that the remained part matters and the removal is almost useless. In Fig.2, we show the original image, 10%-cropped image, and the weight map of the original one. We can observe that even for the block with a large weight, random-crop removes 90% pixels of it. To make full use of the original image and obtain data more reasonably, we introduce compressive sensing.
@@ 3.1 Flexible Sampling Module
FSM is employed to compress the image using compressive sensing with sampling matrix $\Phi$, which is learnable. Due to the high computational complexity, previous sampling applications are limited in datasets with low-resolution images like MNIST and CIFAR-10. To extend the model’s applicability to real-world images, we use block-based compressive sensing algorithm (BCS). BCS is able to compress the image block-by-block. As shown in Fig.3, given an image of size $H\times W$ , we split the image into $L=\frac{H}{B}\times\frac{W}{B}$ non-overlapping blocks, with each block sized $B\times B$. To get the measurements, we have a base sampling matrix $\Phi$ sized $B^{2}\times B^{2}$. Noting $\gamma$ as the sampling ratio, we truncate the first $\gamma B^{2}$ rows of $\Phi$ to get $\Phi_{\gamma}$. Finally, to obtain the measurements, the process can be described as follows:
@@
where $x_{i}\in\mathbf{R}^{B^{2}}$ represents the flattened block of original image, $y_{i}\in\mathbf{R}^{\gamma B^{2}}$ denotes the corresponding measurement, and $i\in[1,L]$. Throughout the whole process, the only parameter is $\Phi$. To accelerate the computation, we can treat every row of $\Phi_{\gamma}$ as a convolution kernel, which turns matrix multiplication into convolution operation. In each training epoch, by randomly selecting CS ratio, we can get different $\Phi_{\gamma}$ and make it closer to the optimal sampling matrix under ratio $\gamma$. Thus, once trained, we can handle all kinds of arbitrary ratios by only one $\Phi$.
Fig. 5: Flowchart of ViT and multi-head self-attention.
Fig. 6: The dual-branch (DB) structure. Each branch contains two fully connected layers for the score and weight prediction.
Table 1: Comparisons of S-IQA v.s. state-of-the-art NR-IQA algorithms on four standard datasets. The best one is shown in red and the second best in blue.
@@
PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
@@
S-IQA-10 0.944 0.936 0.971 0.965 0.930 0.910 0.938 0.933
S-IQA-20 0.975 0.970 0.978 0.970 0.947 0.925 0.942 0.937
S-IQA-50 0.984 0.979 0.983 0.975 0.954 0.944 0.948 0.943
3.2 ViT with Adaptive Embedding Module
Considering that IQA demands a global perception of the whole image, we employ transformer for its excellent performance in constructing long-range correlations. As shown in Fig.4, similar to pure transformer, ViT embeds the input image into the size of $L\times d$, where $d$ denotes the embedding dimension. However, due to the arbitrary CS ratio, the input sequence size $L\times\gamma B^{2}$ is also variable according to the random $\gamma$. To overcome the limitation of the fixed embedding module in ViT, we employ the Adaptive Embedding Module (AEM) to process the measurements. Similar to FSM, we have a base embedding matrix M sized $B^{2}\times B^{2}$. Then, according to the given ratio $\gamma$, we use the first $\gamma B^{2}$ columns of $M$ to form $M_{\gamma}$. Finally, for every measurement $y_{i}$, the embedding process is as follow:
$t_{i}=M_{\gamma}\cdot y_{i}$ (2)
Where $t_{i}\in\mathbf{R}^{d}$ represents the sequence that can be handled by ViT. Noting $T=[t_{1},t_{2},…,t_{L}]$ and $P$ as the positional embedding result, the initial input of ViT is constructed by the following formulation:
$X^{0}=T+P$ (3)
As illustrated in Fig.5, The standard ViT has the same structure as the transformer encoder, containing a stack of $N$ consecutive blocks. Within each block, the attention mechanism is employed to compute correlations between image patches, finally giving high-level features. The computation process is as follows:
$\tilde{X}^{i}$ $=$ $Norm(MSA(X^{i-1})+X^{i-1})$ (4)
$X^{i}$ $=$ $Norm(FF(\tilde{X}^{i})+\tilde{X}^{i})$ (5)
3.3 Dual Branch
$score=\frac{Sum(S(F)\times W(F))}{Sum(W(F))}$ (6)
where $S$ indicates scoring branch and $W$ is weighting branch.
4 EXPERIMENTS
4.1 Experimental Settings
We implement our experiments on two NVIDIA GeForce GTX 1080Ti and a computation platform with PyTorch 1.10.0 and CUDA 11.3 version for the whole training, validating, and testing process. We train and evaluate the performance of our proposed model on four datasets: LIVE[20], CSIQ[21], TID2013[22], KADID-10K[23]. Table 2 shows the summary of the datasets that are used in our experiments.
We choose ViT-B/16-224 as our backbone which is pre-trained on ImageNet-21k with patch size set to 16 and image size 224. This base version transformer contains $N$=12 transformer blocks, with the number of heads being 12 in each layer and the dimension of the feature embedding is set to 768. To be consistent with the patch size used in transformer and protect local information correlations, we set the block size in the sampling model also to 16.
Table 2: Summary of IQA datasets.
Dataset # of DIst. Images # of Dist Types
LIVE 779 5
CSIQ 866 6
TID2013 3000 24
KADID-10K 10,125 25
In the experiments, we randomly split the dataset into 6:2:2 for training, validating, and testing. During training, we set the batch size to 8. We utilize Adam as the optimizer with learning rate $1\times 10^{-5}$ and weight decay $1\times 10^{-5}$. The training loss we use is Mean Square Error (MSE) loss. During testing, we select the one that performs the best in the validation dataset. For each image in the test dataset, we randomly crop a $224\times 224$ sized patch of it for 5 times and obtain the final score by averaging the results. In our experiments, we evaluate our S-IQA with both fixed CS ratios and arbitrary ratios, and several variants of our model are defined with different names. For instance, “S-IQA-10” represents the variant with a fixed CS ratio of 10%, and “S-IQA-r” denotes the variant that deals with an arbitrary CS ratio.
4.2 Evaluation Criteria
Following the prior works, we employ two commonly used criteria: Spearman’s rank-order correlation coefficient (SRCC) and Pearson’s linear correlation coefficient (PLCC). The SRCC is defined as follows:
$SRCC=1-\frac{6\sum_{i=1}^{N}(v_{i}-p_{i})}{N(N^{2}-1)}$ (7)
where $v_{i}$ and $p_{i}$ represent the order of the same image in the ground-truth list and prediction list, and $N$ is the number of images. The PLCC is formulated as:
$PLCC=\frac{\sum_{i=1}^{N}(y_{i}-\mu_{y_{i}})(\tilde{y}_{i}-\mu_{\tilde{y}_{i}})}{\sqrt{\sum_{i=1}^{N}(y_{i}-\mu_{y_{i}})^{2}}\sqrt{\sum_{i=1}^{N}(\tilde{y}_{i}-\mu_{\tilde{y}_{i}})^{2}}}$ (8)
where $y_{i}$ and $\tilde{y}_{i}$ indicates the ground-truth and prediction quality score of $i$-th image, and $\mu_{y_{i}}$ and $\mu_{\tilde{y}_{i}}$ denotes the mean of them. Both of the metrics are in [-1, 1]. In our experiments, a higher value means better performance.
Table 3: In the testing stage, ratio-fixed models perform best under the training CS ratio, while S-IQA-r is rather stable.
@@ 4.2 Evaluation Criteria
(a) PLCC results.
(b) SRCC results.
Fig. 7: PLCC and SRCC of four variants under different CS ratios in CSIQ.
4.3 Comparison with State-of-the-Art Methods
Table 1 shows the overall performance of S-IQA with fixed sampling ratio (10%, 20%, and 50%) on four standard datasets in terms of PLCC and SRCC. We can observe that compared with the methods that utilize the whole image to predict, our proposed S-IQA achieves state-of-the-art performance with much less data. Even if only 10% measurements are available in CSIQ, we can obtain excellent outcomes. Furthermore, one can see that for the same dataset, the performance is getting better with the CS ratio growing, which proves that our sampling module indeed gives more information using a larger CS ratio.
To show the superiority of our FSM, we compare four variants (S-IQA-10/20/50/r) on CSIQ with the CS ratio ranging from 0.01 to 1.0. As shown in Fig.7, except for S-IQA-r, every model with fixed ratio tends to perform better as the CS ratio grows from 0.01. While after the peak, the performance falls fast for the reason that only the first few rows of the sampling matrix are trained (e.g., the first 10% rows of the matrix in S-IQA-10), and as the ratio becomes larger, more and more untrained (i.e., random) parameters participate in calculating and finally destroy the result. In contrast, the whole matrix of FSM in S-IQA-r is trained so it can deal with any ratio once trained. Moreover, the results of S-IQA-r are pretty stable when the CS ratio varies. But Table 3 indicates that despite the flexibility, S-IQA-r cannot beat ratio-fixed variants when the testing ratio exactly equals that of training.
4.4 Visualization
To figure out exactly how the dual-branch structure helps when predicting the quality score, we generate different feature maps using S-IQA-10. As shown in Fig.8, we add White Gaussian Noise to the right half of the image according to different SNRs (10, 1, and 0.1). Similar to what we have assumed, different patches of the image matter differently and the map tends to give more weight to the outline or background. We can observe that with more severe distortion, the score and weight map both give lower values. Among the two, the weight map is dominant and is easier to be affected by noise.
Table 4: Ablation results of ViT and DB at CS ratio=10%.
@@ 4.4 Visualization
[…]
Fig. 8: Dual branch feature maps of an image with different degrees of distortion on the right half. The first row is the pristine image and the SNR from the second row down is 10, 1, and 0.1.
4.5 Ablation Study
In this section, we conduct ablation study on S-IQA-10 to prove the effectiveness of the transformer backbone and the dual branch. For the transformer part, we use ResNet-101 to replace it for feature extraction. In the research of dual branch (DB), we replace it with two layers of CNN and MLP. The result is shown in Table 4. We can observe that each part has a certain effect on the quality prediction. Transformer has the superiority of constructing long-range correlations and the dual-branch can reasonably allocate weight among the patches.
5 CONCLUSION
In this paper, we propose a new data-efficient framework for Image Quality Assessment (IQA) dubbed S-IQA. It first samples the image using the FSM to obtain measurements at any given CS ratio. Then we utilize the AEM to overcome the problem that measurements don’t match the fixed input shape of the transformer. After embedding, the measurements are fed into the Vision Transformer for further feature extraction. Finally, the DB weights every patch of the image and predicts the final quality score. Experimental results demonstrate that with less data usage, our proposed S-IQA outperforms many recent state-of-the-art NR-IQA methods.

### v2

# Image Quality Assessment With Compressed Sampling

@@ Abstract
No-Reference Image Quality Assessment (NR-IQA) aims at estimating image quality in accordance with subjective human perception. However, most methods focus on exploring increasingly complex networks to improve the final performance, accompanied by limitations on input images. Especially when applied to high-resolution (HR) images, these methods offen have to adjust the size of original image to meet model input. To further alleviate the aforementioned issue, we propose two networks for NR-IQA with Compressive Sampling (dubbed CL-IQA and CS-IQA). They consist of four components: (1) The Compressed Sampling Module (CSM) to sample the image (2) The Adaptive Embedding Module (AEM). The measurements are embedded by AEM to extract high-level features. (3) The Vision Transformer and Scale Swin TranBlocksformer Moudle (SSTM) to extract deep features. (4) The Dual Branch (DB) to get final quality score. Experiments show that our proposed methods outperform other methods on various datasets with less data usage.
Index Terms:
Image Quality Assessment, Compressive Sensing, Vision Transformer.
I Introduction
While utilizing images from the internet, we should be mindful of their potential transformation journey. These images might have been captured by a camera, compressed for sharing online and disseminated widely before finally reaching us. These low-quality images can affect the our visual feelings and even cause deadly problems, especially in autonomous driving. For the reason, it is crucial to predict the perceptual image quality in our daily lives.
@@ I Introduction
Objective Image Quality Assessment (IQA) is an approach that uses computational models to ascertain the perceived quality of an image from a human perspective. The objective quality metrics can be categorized into two based on whether a lossless reference image is available: full-reference (FR-IQA)[1, 2, 3, 4] and no-reference (NR-IQA). Unlike FR-IQA, NR-IQA[5, 6, 7] does not have access to reference images which is harder and more widely used. With the success of deep learning and convolution neural networks (CNNs) in computer vision tasks, CNN-based methods have significantly outperformed traditional approaches in handling real-world distortions. Early deep learning NR-IQA methods employed stacked CNNs for feature extraction. Ma[8] proposes a multi-task network where two sub-networks are trained for distortion identification and quality assessment. Hyper-IQA[9] leverages both low-level and high-level features and makes the latter redirect the former. Furthermore, Zhu[10] proposes a model that employs meta-learning to capture shared prior knowledge among different distortions. But in the era of high-resolution Image Quality Assessment, both methods are faced with substantial computational challenges while processing entire images. There already exist some data-efficient methods for quality assessment, such as FAST-VQA[11] with random-crop. But cropping cannot guarantee that the remained part matters.
Compressive sensing is a novel method that breaks through the limitations of the Nyquist theorem on signal sampling[12, 13]. This technique shows that one can faithfully reconstruct an entire image from a minimal amount of measurements compared to the original image, thus significantly simplifying the sampling process.
The concept of compressive learning (CL) was first proposed by Calderbank et al.[14] and Davenport et al.[15] when they built the inference system directly using measurements without reconstruction. Tran[16] makes some theoretical works and gradually finds some multidimensional properties of CL. A recent work [17] by Chong Mou et al. employs transformer as the backbone to cope with information loss through element-wise correlations and achieves amazing performance.
In many applications, such as NR-IQA, the precise reconstructed image isn’t a primary concern. Still, we emphasize more on the final score, which means we can directly perform high-level tasks after initialization or even on the compressed domain.Inspired by this work, to make NR-IQA more data-efficient and extract the feature of the original image globally, we introduce compressive sensing in NR-IQA task.We propose CL-IQA that integrates compressive learning with the NR-IQA task and CS-IQA which skip the AEM block of CL-IQA. They share the same backbone. Firstly, CSM allow us obtain measurements at an arbitrary ratio. Secondly, the measurements are adaptive embedded by AEM. Thirdly, the measurements or embedded measurements were fed into the vision transformer and SSTM to extract deep features. Finally, a dual-branch structure is employed for the score prediction. Our main contributions are as follows:
Figure 1: Random-crop removes pixels from every block despite the importance, leading to severe information loss.
$\bullet$ We apply compressive learning to image quality assessment with, enable sampling of images at arbitrary ratios. To our knowledge, CL-IQA is the first work integrating compressive learning and image quality assessment.
$\bullet$ We pretrain the CSM with lightweight compressive sensing network enable it captures the feature about quality assessment more effectively. The proposed CS-IQA significantly reduce the complexity by skip the AEM module.
$\bullet$ Experiments show that CL-IQA and CS-IQA outperforms state-of-the-art IQA methods with benefiting from the sampling modules.
Figure 2: The architecture of our proposed CS-IQA and CL-IQA.
II Proposed Approach
@@
In this section, we elaborate on the details of our framework and integrated model. The whole architecture is presented in Fig.2, which is mainly composed of a Compressed Sampling Module (CSM), a Adaptive Embedding Module (AEM), a Vision Transformer, a Scale Swin TranBlocksformer Moudle (SSTM) and a Dual Branch (DB).
II-A Compressed Sampling Module
In Fig.1, we show the original image, 10%-cropped image, and the weight map of the original one. We can observe that even for the block with a large weight, random-crop removes 90% pixels of it. Actually, instead of the remove pixels randomly, we might only need a small number of measurements sampled globally from original image (perhaps 25% measurements or even less) to achieve highly reliable score prediction. In implementation, we employ compressive sensing for sampling.
@@ II-A Compressed Sampling Module
CSM is employed to compress the image using compressive sensing with sampling matrix $\Phi$, which is learnable. Due to the high computational complexity, previous sampling applications are limited in datasets with low-resolution images. To extend the model’s applicability to real-world images, we use block-based compressive sensing algorithm (BCS)[18]. BCS is able to compress the image block-by-block. As shown in Fig.2, given an image of size $H\times W$, we split the image into $L=\frac{H}{B}\times\frac{W}{B}$ non-overlapping blocks, with each block sized $B\times B$. To get the measurements, we have a base sampling matrix $\Phi$ sized $B^{2}\times B^{2}$. Noting $\gamma$ as the sampling ratio, we truncate the first $\gamma B^{2}$ rows of $\Phi$ to get $\Phi_{\gamma}$. Finally, to obtain the measurements, the process can be described as follows:
@@
where $x_{i}\in\mathbf{R}^{B^{2}}$ represents the flattened block of original image, $y_{i}\in\mathbf{R}^{\gamma B^{2}}$ denotes the corresponding measurement, and $i\in[1,L]$. Throughout the whole process, the only parameter is $\Phi$. To accelerate the computation, we can treat every row of $\Phi_{\gamma}$ as a convolution kernel, which turns matrix multiplication into convolution operation.
There are many CS methods but some of them have extremely high computational complexity due to the iterative calculations, such as ISTA-Net[19],AMS-Net[20] ,OPINE-Net[21]. Shi et al.[22] propose a lightweight network model called CSNet and show its promising performance. Therefore we pretrain our CSM with CSNet.
Figure 3: Process of Flexible Sampling Module (AEM).
II-B Adaptive Embedding Module
Considering that IQA demands a global perception of the whole image, we employ transformer for its excellent performance in constructing long-range correlations. Similar to pure transformer, ViT embeds the input image into the size of $L\times d$, where $d$ denotes the embedding dimension. However, due to the arbitrary CS ratio, the input sequence size $L\times\gamma B^{2}$ of CL-IQA is also variable according to the parameter $\gamma$. As shown in Fig.3, to overcome the limitation of the fixed embedding module in ViT, we have a base embedding matrix M sized $B^{2}\times B^{2}$. Then, according to the given ratio $\gamma$, we use the first $\gamma B^{2}$ columns of $M$ to form $M_{\gamma}$. Finally, for every measurement $y_{i}$, the embedding process is as follow:
$t_{i}=M_{\gamma}\cdot y_{i}$ (2)
Where $t_{i}\in\mathbf{R}^{d}$ represents the sequence that can be handled by ViT.
II-C Vision Transformer
Noting $T=[t_{1},t_{2},…,t_{L}]$ and $P$ as the positional embedding result, the initial input of ViT is constructed by the following formulation:
$X^{0}=T+P$ (3)
The standard ViT[23] has the same structure as the transformer encoder, containing a stack of $N$ consecutive blocks. Within each block, the attention mechanism is employed to compute correlations between image patches, finally giving high-level features. The computation process is as follows:
$\tilde{X}^{i}$ $=$ $Norm(MSA(X^{i-1})+X^{i-1})$ (4)
$X^{i}$ $=$ $Norm(FF(\tilde{X}^{i})+\tilde{X}^{i})$ (5)
II-D Scale Swin TranBlocksformer Moudle
The Scale Swin Transformer Moudle consists of Swin Transformer Layers[24] and a convolutional layer. The SSTM first encodes the input feature through 2 layers of STL:
$X_{1}^{i}=S_{STL}(X_{0}^{i})$ (6)
Then the convolutional layers are applied before the residual connection. The output of SSTB is formulated as:
$X_{1}^{i}=\alpha\cdot S_{CONV}(X_{1}^{i})+X_{1}^{i}$ (7)
the coefficient $\alpha$ denotes the scale factor of the output of 2 layers of STL.
Figure 4: The dual-branch (DB) structure. Each branch contains two fully connected layers for the score and weight prediction.
II-E Dual Branch
$score=\frac{Sum(S(F)\times W(F))}{Sum(W(F))}$ (8)
where $S$ indicates scoring branch and $W$ is weighting branch.
TABLE I: Comparisons with state-of-the-art NR-IQA algorithms.
@@ II-E Dual Branch
Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
@@
CL-IQA-10 0.944 0.936 0.971 0.965 0.930 0.910 0.938 0.933
CL-IQA-20 0.975 0.970 0.978 0.970 0.947 0.925 0.942 0.937
CL-IQA-50 0.984 0.979 0.983 0.975 0.954 0.944 0.948 0.943
CS-IQA 0.947 0.944 0.973 0.965 0.940 0.926 0.945 0.943
TABLE II: ratio-fixed models perform best under the training CS ratio, while CL-IQA-r is rather stable.
@@
TABLE III: Ablation results of ViT and DB at CS ratio=10%.
@@
0.901/0.890 0.946/0.930 0.897/0.868 0.917/0.911
0.938/0.934 0.966/0.952 0.921/0.897 0.933/0.926
0.944/0.936 0.971/0.965 0.930/0.910 0.938/0.933
III Experiments
III-A Experimental Settings
We implement our experiments on NVIDIA GeForce GTX 1080Ti with PyTorch 1.10.0 and CUDA 11.3 version for the whole training, validating, and testing process. We train and evaluate the performance of our proposed model on four datasets: LIVE[32], CSIQ[33], TID2013[34], KADID-10K[35]. And T91 dataset is used to pretrain the CSM.
We choose ViT-B/8-224[23] as our backbone which is pre-trained on ImageNet-21k with patch size set to 8 and image size 224. This base version transformer contains $N$=12 transformer blocks, with the number of heads being 12 in each layer and the dimension of the feature embedding is set to 768. To be consistent with the patch size used in transformer and protect local information correlations, we set the block size in the sampling model also to 16.
In the experiments, we randomly split the dataset into 8:2 for training and testing. During training, we set the batch size to 8. We utilize Adam as the optimizer with learning rate $1\times 10^{-5}$ and weight decay $1\times 10^{-5}$. The training loss we use is Mean Square Error (MSE) loss. During testing, we select the one that performs the best in the validation dataset. In CL-IQA, for each image in the test dataset, we randomly crop a $224\times 224$ sized patch of it for 5 times and obtain the final score by averaging the results. In our experiments, we evaluate our CL-IQA with both fixed CS ratios and arbitrary ratios. For instance, “CL-IQA-10” represents the variant with a fixed CS ratio of 10%, and “CL-IQA-r” denotes the variant that deals with an arbitrary CS ratio.
III-B Comparison with State-of-the-Art Methods
Table I shows the overall performance of CL-IQA with fixed sampling ratio (10%, 20%, and 50%) and CS-IQA on four standard datasets in terms of PLCC and SRCC. We can observe that compared with the methods that utilize the whole image to predict, our proposed methods achieves state-of-the-art performance with much less data. Even if only 10% measurements are available in CSIQ, we can obtain excellent outcomes. Furthermore, one can see that for the same dataset, the performance is getting better with the CS ratio growing.
Figure 5: The dual-branch (DB) structure. Each branch contains two fully connected layers for the score and weight prediction.
III-C Visualization
To figure out exactly how the dual-branch structure helps when predicting the quality score, we generate different feature maps using CL-IQA-10. As shown in Fig.5, we add White Gaussian Noise to the right half of the image according to different SNRs(Signal-to-Noise Radio) (10, 1, and 0.1). As what we have assumed, different patches of the image matter differently and the map tends to give more weight to the outline or background.
III-D Ablation Study
In this section, we conduct ablation study on CL-IQA-10 to prove the effectiveness of the transformer backbone and the dual branch. For the transformer part, we use ResNet-101 to replace it for feature extraction. In the research of dual branch (DB), we replace it with two layers of CNN and MLP. The result is shown in Table II. We can observe that each part has a certain effect on the quality prediction. Transformer has the superiority of constructing long-range correlations and the dual-branch can reasonably allocate weight among the patches.
CONCLUSION
In this paper, we propose a new data-efficient framework for NR-IQA dubbed CL-IQA and CS-IQA. It samples the image using the Compressed Sampling Module to obtain measurements globally. Then we utilize the adaptive embedding to overcome the problem that measurements don’t match the fixed input shape of the transformer. The measurements are fed into the Vision Transformer and Scale Swin TranBlocksformer Moudle for further feature extraction. Finally, the DB weights every patch of the image and predicts the final quality score. Experiments demonstrate that our proposed methods outperforms many recent state-of-the-art NR-IQA methods.